# SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation of Large Language Models

### v2

@@ 8.2 Ethical Concerns
The PEFT method, while efficient, raises critical ethical concerns regarding the security of built-in safety measures in AI models. As demonstrated in (Lermen & Rogers-Smith, 2024), subversive fine-tuning techniques can bypass safety training intended to prevent the generation of harmful content. The ease and affordability of such methods underscore the vulnerability of safety protocols. It is imperative to develop robust safeguards that keep pace with technological advancements, ensuring that efficiency gains in model tuning do not compromise the ethical use of AI.
@@ 9 Acknowledgment
We would like to extend our gratitude to Prof. Jianyong Wang from Tsinghua University and the Computer Science Department of the Advanced Project Research Laboratory of Tsinghua University High School for providing guidance to the research.
Appendix A Training Details
@@ A.1 Analysis
For the singular values and vectors analysis in section 3, we apply partial fine-tuning, LoRA SORSA (with and without orthonormal regularizer) on Llama 2 7B (Touvron et al., 2023) model, using the same training data and optimizer as A.2. In the analysis, LoRA and SORSA are only applied to q_proj and v_proj matrices, respectively. FT is only trainable on q_proj and v_proj. We only calculate the loss on the response part. The models are trained with BF16 mix precision. See Table 2 for hyperparameters.
@@ A.2 NLG Experiments
For our NLG tasks, we adapt Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by LoRA (Hu et al., 2021), PiSSA (Meng et al., 2024) and SORSA, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset, and test by GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) dataset, We use AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. SORSA adapters are applied on all linear matrices in every layer. We only calculate the loss on the response part. The models are trained with BF16 mix precision. See Table 3 for hyperparameters.

### v3

@@ 9.2 Ethical Concerns
The PEFT method, while efficient, raises critical ethical concerns regarding the security of built-in safety measures in AI models. As demonstrated in Lermen & Rogers-Smith (2024), subversive fine-tuning techniques can bypass safety training intended to prevent the generation of harmful content. The ease and affordability of such methods underscore the vulnerability of safety protocols. It is imperative to develop robust safeguards that keep pace with technological advancements, ensuring that efficiency gains in model tuning do not compromise the ethical use of AI.
@@ 10 Acknowledgment
We would like to extend our gratitude to Prof. Jianyong Wang from Tsinghua University and the Computer Science Department of the Advanced Project Research Laboratory of Tsinghua University High School for providing the endorsement to the research.
Appendix A Experiments Details
@@ A.1 Analysis
For the singular values and vectors analysis in Section 3, we applied partial fine-tuning, LoRA and SORSA (with and without orthonormal regularizer) on Llama 2 7B (Touvron et al., 2023) model, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset. In the analysis, LoRA and SORSA were only applied to q_proj and v_proj matrices, respectively. For Partial FT, we set model’s q_proj and v_proj matrices to trainable. We only calculated the loss on the response part. The models are trained with TF32 + BF16 mix precision. See Table 2 for hyperparameters.
We also found we should only perform SVD for analysis using CPU, in order to get the precise analysis data.
@@ A.2 NLG Experiments
For our NLG tasks, we adapted Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by SORSA. For GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) evaluations, we trained those models with the first 100K data in MetaMathQA (Yu et al., 2024) dataset. For HumanEval (Chen et al., 2021) evaluation, we use the first 100K data in CodeFeedback Filtered Instruction(Zheng et al., 2024) dataset.
We used AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. SORSA adapters were applied on all linear matrices in every layer. We only calculated the loss on the response part. The models are trained with TF32 + BF16 mix precision. See Table 3 for hyperparameters.
@@
1@@ Instruction
2Here is the given code to do completion:
3‘‘‘python
4{prompt}
5‘‘‘
6
7Please continue to complete the function with python programming language. You are not allowed to modify the given code and do the completion only.
8
9Please return all completed codes in one code block.
10This code block should be in the following format:
11’’’python
12# Your codes here
13’’’
14
15@@ Response
Listing 1: Prompt used for HumanEval evaluation.
Appendix B Proofs
B.1 Proof of Theorem 5.1
Proof.
We prove this in two steps:
First, we show that $f(U_{p})=\|U^{\top}_{p}U_{p}-I\|_{F}$ is convex. Then, we prove that $g(V_{p})=\|V_{p}V^{\top}_{p}-I\|_{F}$ is convex.
Since the sum of convex functions is convex, this will establish the convexity of ${\mathcal{L}}_{reg}$.
Let $U_{p},W\in\mathbb{R}^{m\times r}$. The Hessian of $f$ at $U_{p}$ in the direction $W$ is given by:
$\begin{split}\nabla^{2}f(U_{p})[W,W]&=\lim_{\epsilon\to 0}\frac{1}{\epsilon^{2}}\Big(f(U_{p}+\epsilon W)-2f(U_{p})+f(U_{p}-\epsilon W)\Big)\\ &=\lim_{\epsilon\to 0}\frac{1}{\epsilon^{2}}\Big(\|(U_{p}+\epsilon W)^{\top}(U_{p}+\epsilon W)-I\|_{F}\\ &\quad-2\|U^{\top}_{p}U_{p}-I\|_{F}+\|(U_{p}-\epsilon W)^{\top}(U_{p}-\epsilon W)-I\|_{F}\Big)\\ &=\lim_{\epsilon\to 0}\frac{1}{\epsilon^{2}}\Big(\|U^{\top}_{p}U_{p}+\epsilon(U^{\top}_{p}W+W^{\top}U_{p})+\epsilon^{2}W^{\top}W-I\|_{F}\\ &\quad-2\|U^{\top}_{p}U_{p}-I\|_{F}+\|U^{\top}_{p}U_{p}-\epsilon(U^{\top}_{p}W+W^{\top}U_{p})+\epsilon^{2}W^{\top}W-I\|_{F}\Big)\\ &=2\|W^{\top}W\|_{F}\end{split}$ (14)
Since $\|W^{\top}W\|_{F}\geq 0$ for all $W$, we have $\nabla^{2}f(U_{p})[W,W]\geq 0$, which proves that $f$ is convex.
The proof for $g(V_{p})$ follows the same steps as for $f(U_{p})$, leading to the same conclusion.
Therefore, both $f(U_{p})$ and $g(V_{p})$ are convex, and consequently, ${\mathcal{L}}_{reg}$ is convex. ∎
B.2 Proof of Theorem 5.2
Proof.
To prove Lipschitz continuity, we need to show that there exists a constant $L>0$ such that for any two pairs of matrices $(U_{1},V_{1})$ and $(U_{2},V_{2})$:
$|{\mathcal{L}}_{reg}(U_{1},V_{1})-{\mathcal{L}}_{reg}(U_{2},V_{2})|\leq L(\|U_{1}-U_{2}\|_{F}+\|V_{1}-V_{2}\|_{F})$ (15)
Let’s break this down into two parts:
First, consider $|\|U_{1}^{\top}U_{1}-I\|_{F}-\|U_{2}^{\top}U_{2}-I\|_{F}|$:
$\begin{split}|\|U_{1}^{\top}U_{1}-I\|_{F}-\|U_{2}^{\top}U_{2}-I\|_{F}|&\leq\|(U_{1}^{\top}U_{1}-I)-(U_{2}^{\top}U_{2}-I)\|_{F}\\ &=\|U_{1}^{\top}U_{1}-U_{2}^{\top}U_{2}\|_{F}\\ &=\|U_{1}^{\top}U_{1}-U_{1}^{\top}U_{2}+U_{1}^{\top}U_{2}-U_{2}^{\top}U_{2}\|_{F}\\ &\leq\|U_{1}^{\top}(U_{1}-U_{2})\|_{F}+\|(U_{1}^{\top}-U_{2}^{\top})U_{2}\|_{F}\\ &\leq\|U_{1}^{\top}\|_{F}\|U_{1}-U_{2}\|_{F}+\|U_{1}-U_{2}\|_{F}\|U_{2}\|_{F}\\ &\leq(\|U_{1}\|_{F}+\|U_{2}\|_{F})\|U_{1}-U_{2}\|_{F}\end{split}$ (16)
Here, we’ve used the triangle inequality and the sub-multiplicative property of the Frobenius norm.
Similarly for $V$:
$|\|V_{1}^{\top}V_{1}-I\|_{F}-\|V_{2}^{\top}V_{2}-I\|_{F}|\leq(\|V_{1}\|_{F}+\|V_{2}\|_{F})\|V_{1}-V_{2}\|_{F}$ (17)
Combining these results:
$\begin{split}|{\mathcal{L}}_{reg}(U_{1},V_{1})-{\mathcal{L}}_{reg}(U_{2},V_{2})|&\leq(\|U_{1}\|_{F}+\|U_{2}\|_{F})\|U_{1}-U_{2}\|_{F}\\ &\quad+(\|V_{1}\|_{F}+\|V_{2}\|_{F})\|V_{1}-V_{2}\|_{F}\\ &\leq\max(\|U_{1}\|_{F}+\|U_{2}\|_{F},\|V_{1}\|_{F}+\|V_{2}\|_{F})\\ &\quad\cdot(\|U_{1}-U_{2}\|_{F}+\|V_{1}-V_{2}\|_{F})\end{split}$ (18)
Let $L_{reg}=\max(\|U_{1}\|_{F}+\|U_{2}\|_{F},\|V_{1}\|_{F}+\|V_{2}\|_{F})$. This $L$ is finite because the Frobenius norms of $U$ and $V$ are bounded (they represent orthonormal matrices in the ideal case).
Therefore, we have shown that:
$|{\mathcal{L}}_{reg}(U_{1},V_{1})-{\mathcal{L}}_{reg}(U_{2},V_{2})|\leq L_{reg}(\|U_{1}-U_{2}\|_{F}+\|V_{1}-V_{2}\|_{F})$ (19)
This proves that ${\mathcal{L}}_{reg}$ is Lipschitz continuous with Lipschitz constant $L_{reg}$. ∎
B.3 Proof of Theorem 5.3
Proof.
For gradient descent to converge while the $\eta_{t}=\eta_{d}$, we need:
$\eta_{d}<\frac{2}{L}=\frac{2}{L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}}$ (20)
Rearranging this inequality:
$\begin{split}\eta_{d}(L_{train}+\frac{\gamma}{\eta_{d}}L_{reg})&<2\\ \eta_{d}L_{train}+\gamma L_{reg}&<2\\ \gamma L_{reg}&<2-\eta_{d}L_{train}\\ \gamma&<\frac{2-\eta_{d}L_{train}}{L_{reg}}\end{split}$ (21)
$0\leq\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (22)
This establishes the upper bound for $\gamma$. ∎
B.4 Proof of Lemma 5.4
Proof.
First, let’s consider the effect of the orthonormal regularizer. The regularizer aims to make $U_{p}^{\top}U_{p}\approx I$ and $V_{p}V_{p}^{\top}\approx I$. We can quantify this approximation as:
$\nabla_{W_{p}}{\mathcal{L}}_{reg}\leq\epsilon_{\nabla}$ (23)
where $\epsilon_{\nabla}>0$ is a small constant.
Now, let’s consider the difference between $W_{p}^{\text{unreg}}$ and $W_{p}^{\text{reg}}$:
$\|W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}\|_{F}=\gamma\|\nabla_{W_{p}}{\mathcal{L}}_{reg}\|_{F}$ (24)
Now, we can use Weyl’s inequality, which states that for matrices A and B:
$|\sigma_{i}(A+B)-\sigma_{i}(A)|\leq\|B\|_{2}\leq\|B\|_{F}$ (25)
Applying this to our case, with $A=W_{p}^{\text{unreg}}$ and $B=W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}$:
$|\sigma_{i}^{\text{reg}}-\sigma_{i}^{\text{unreg}}|\leq\|W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}\|_{F}\leq\gamma\epsilon_{\nabla}$ (26)
Let $\epsilon=\gamma\epsilon_{\nabla}$. Then we have:
$-\epsilon\leq\sigma_{i}^{\text{reg}}-\sigma_{i}^{\text{unreg}}\leq\epsilon$ (27)
Rearranging this inequality gives us our desired bound:
$(1-\epsilon)\sigma_{i}^{\text{unreg}}\leq\sigma_{i}^{\text{reg}}\leq(1+\epsilon)\sigma_{i}^{\text{unreg}}$ (28)
B.5 Proof of Theorem 5.5
Proof.
Let $W_{p}=U_{p}\Sigma_{p}V_{p}^{\top}$ be the principal part of the singular value decomposition approximation of $W$. The condition number is given by:
$\kappa(W_{p})=\frac{\sigma_{\max}(W_{p})}{\sigma_{\min}(W_{p})}$ (29)
where $\sigma_{\max}$ and $\sigma_{\min}$ are the maximum and minimum singular values of $W_{p}$.
At initialization ($t=0$): Due to SVD initialization, $U_{p}$ and $V_{p}^{\top}$ are perfectly orthonormal, so:
$\kappa(U_{p}^{\text{unreg}})=\kappa((V_{p}^{\text{unreg}})^{\top})=\kappa(U_{p}^{\text{reg}})=\kappa((V_{p}^{\text{reg}})^{\top})=1$ (30)
And $\epsilon_{0}=0$, $\delta_{1,0}=\delta_{2,0}=0$. Therefore:
$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=\frac{\kappa(\Sigma_{p}^{\text{reg}})}{\kappa(\Sigma_{p}^{\text{unreg}})}=1$ (31)
During training ($t>0$): As training progresses, $U_{p}$ and $V_{p}^{\top}$ deviate from orthonormality in the unregularized case. We quantify this deviation:
$\|U_{p}^{\top}U_{p}-I\|_{F}\leq\epsilon_{1,t}$ (32)
$\|V_{p}^{\top}V_{p}-I\|_{F}\leq\epsilon_{2,t}$ (33)
where $\epsilon_{1,t},\epsilon_{2,t}>0$ increase over time t.
For the regularized matrices, we can bound their condition numbers:
$\kappa(U_{p}^{\text{reg}})\leq 1+\delta_{1,t}$ (34)
$\kappa(V_{p}^{\text{reg}})\leq 1+\delta_{2,t}$ (35)
where $\delta_{1,t},\delta_{2,t}$ are small positive numbers that remain bounded due to the regularization.
From the Lemma 5.4, we arrive at:
$\begin{split}\frac{(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1-\epsilon_{t}}{1+\epsilon_{t}})}{\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})}\leq\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}\leq\frac{(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})}{\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})}\end{split}$ (36)
As training continues, in the unregularized case, $\kappa(U_{p,\text{unreg}})$ and $\kappa(V_{p,\text{unreg}}^{\top})$ tend to increase as $U_{p}$ and $V_{p}^{\top}$ deviate further from orthonormality. On the other hand, $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ will approach to 1 because of the reinforcement in orthonormality. Therefore, $\exists c>0$, while $t>c$:
$\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})>(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ (37)
will hold true.
Therefore, while $t>c$, we have:
$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}<1$ (38)
indicating an improvement in the condition number. ∎